A chemist wants to mix a 64% alcohol solution with 7 liters of 51% alcohol solution to produce a solution that is 62% alcohol. How many liters of the 64% alcohol solution should be used? Round to two decimal places.
Answer by lwsshak3(11628)

A chemist wants to mix a 64% alcohol solution with 7 liters of 51% alcohol solution to produce a solution that is 62% alcohol. How many liters of the 64% alcohol solution should be used? Round to two decimal places.
let x=amt of 64% solution to be used
x+7=amt of 62% mixed solution
64%x+51%*7=62%(x+7)
.64x+3.57=.62x+4.34
0.2x=0.77
x=3.85
amt of 64% solution to be used=3.85 liters